OpenStudy (anonymous):

Rotate the axes of to get rid of the cross product term of 9x^2+y^2+6xy=4? I believe I have to find the symmetric coefficient matrix and diagonalize it, but I'm not sure how to do that.

OpenStudy (anonymous):

Well: \[9x^2+y^2+6xy=4 \rightarrow \vec{r}^T M \vec{r}=4 \implies (x,y)\left[\begin{matrix}9 & 3 \\ 3 & 1\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right)=4\]

OpenStudy (anonymous):

So you want to diagnolize M I take it.

OpenStudy (anonymous):

Yes! Is that the correct way to solve the problem?

OpenStudy (anonymous):

Yes, you will need to diagnolize it. This is a better explanation: http://ltcconline.net/greenl/courses/203/MatrixOnVectors/conics.htm

OpenStudy (anonymous):

Thank you!